OpenStudy (anonymous):

Find the equation of the line containing the point P and having the slope m. Write your answer in slope-intercept form. Graph this line and submit the graph to the Dropbox. Please be sure to show all of your work. P = (-4, -6), m = 3

OpenStudy (anonymous):

\[\Large \begin{array}{l} P = ( - 4, - 6)\\ m = 3\\ y = mx + c\\ - 6 = 3( - 4) + c\\ - 6 = - 12 + c\\ c = 6\\ y = 3x + 6 \end{array}\]

OpenStudy (campbell_st):

use the point gradient formula \[y - y _{1} = m(x - x _{1})\] from your information \[y _{1} = -6 x _{1} = - 4\] then \[y - (-6) = 3(x - (-4))\] simplify and put in the form... y = mx + b

OpenStudy (anonymous):

Thank you both for the help.